# EVA-MILP: Towards Standardized Evaluation of MILP Instance Generation

### v2

@@ 6. Performance Evaluation
L.2 Validation Experiment
More validation experiments for solver-internal features are in the supplementary files.
L.3 Some Insights
We noted a stark divergence in the performance of the same G2MILP model on IS and SC (Set Covering) problems. For SC problems, G2MILP largely replicated the performance characteristics of the original instance set, whereas for IS problems, the difficulty of the generated instances increased dramatically. This could be attributed to the following reasons:
For the IS problem, the objective is to find the largest subset of nodes in a graph such that no two nodes within the subset are adjacent [37, 38]. In the MILP formulation, a binary variable $x_{i}$ is typically assigned to each node $i$ (indicating whether node $i$ is in the independent set), and a constraint $x_{i}+x_{j}\leq 1$ is added for each edge $(i,j)$ in the graph. In G2MILP’s bipartite graph representation, “constraint nodes” correspond to these edge constraints. Thus, when G2MILP modifies a “constraint node”, it is altering a constraint associated with a specific edge in the graph. The “structure and hardness” of IS problems are predominantly derived from the global topological properties of the original graph, rather than merely the presence or absence of individual edge constraints or their coefficients (which are typically 1). G2MILP, by modifying individual or a few edge constraints, may struggle to effectively learn or preserve these critical global graph attributes. It might only be performing local “poking” or “connecting” operations, the global structural impact of which can be drastic and unpredictable. The paper mentions that G2MILP “iteratively corrupts and replaces parts of the original graphs”. For a problem like IS, which is sensitive to global structure, such local, iterative replacements might more readily lead to structural deviations.
For the SC problem, constraints directly define the covering requirements [20]. Modifications to constraints by G2MILP directly manipulate the core semantic units of the problem. The “structure and hardness” of SC instances in the training data are primarily manifested in the configuration of these constraints (i.e., which constraints exist and which variables they involve) [20]. Therefore, G2MILP’s Variational Autoencoder (VAE) has a higher probability of learning effective patterns within these configurations.
This further underscores that structural similarity, in itself, is not conclusively informative and proves useful only for certain problems. This motivates the need for our proposed, more general evaluation metric.
Appendix M Limitations
While we advocate for the EVA-MILP framework and its emphasis on solver-internal features as a significant advancement, potential limitations warrant acknowledgment.
  • Metric Selection and Evolution: The initial selection of ’useful metrics,’ though comprehensive, may necessitate ongoing refinement. Future work should consider incorporating emerging or problem-specific indicators to ensure the framework remains current and to avoid potential inherent biases in the chosen metrics.
  • Solver-Specific Feature Dependency: Some internal features are inherently dependent on the specific solver used for evaluation. This requires careful consideration when generalizing findings across different solvers or when comparing instances evaluated with disparate solver technologies. The characteristics observed might, in part, reflect the solver’s architecture rather than solely the instance’s intrinsic properties.
  • Information Granularity from Solvers: A challenge encountered during our work was the level of detail disclosed by some commercial solvers. For instance, when logging the behavior of Gurobi, specific details about the types of heuristics employed were not always available. This lack of transparency can hinder a deeper analysis of instance properties and solver interactions. To mitigate this, we suggest that future research could benefit from utilizing open-source solvers like SCIP [34] or HiGHS [44], which may offer greater access to, and control over, internal algorithmic information. The more comprehensive the information provided by a solver, the more effectively researchers can analyze instance features and hardness.
  • Interpretation Expertise: Although EVA-MILP aims to simplify and standardize the evaluation process, comprehensively interpreting the multi-faceted data generated (spanning mathematical validity, structural similarity, computational hardness, and solver behavior) may still demand considerable expertise in optimization and solver mechanisms. Developing tools or guidelines for more automated or guided interpretation remains an area for future development."
Appendix N Broader Impact
The introduction of the EVA-MILP framework has several potential broader impacts on the fields of mathematical optimization and machine learning:
Positive Impacts:
  • Advancing Research in Combinatorial Optimization: By providing a standardized and comprehensive methodology for evaluating MILP instance generation, EVA-MILP can foster more rigorous and comparable research. This can accelerate the development of higher-quality synthetic instances, which are crucial for testing, benchmarking, and improving MILP solvers.
  • Enhancing Machine Learning for Optimization: The availability of diverse, well-characterized, and challenging MILP instances is vital for training and validating machine learning models aimed at improving optimization algorithms (e.g., for tasks like branching, node selection, or parameter tuning). EVA-MILP can help ensure that the synthetic data used for these purposes more accurately reflects the complexities of real-world problems, leading to more robust and effective ML-driven optimization techniques.
  • Development of More Realistic Benchmarks: The framework encourages a deeper understanding of instance features beyond superficial structural similarity, pushing for the creation of synthetic benchmarks that better capture the computational hardness and nuanced characteristics of operational problems. This can lead to solvers that are better equipped to handle real-world challenges.
  • Facilitating Fairer Comparisons: EVA-MILP offers a more level playing field for comparing different instance generation techniques, moving beyond limited or inconsistent evaluation practices. This transparency can guide researchers and practitioners in selecting or developing generators best suited for their specific needs.
  • Educational Tool: The framework and its associated metrics can serve as an educational resource for students and researchers new to MILP, providing insights into what constitutes a ’good’ or ’hard’ instance and how various features influence solver performance.
Potential Negative Impacts and Mitigations:
  • Over-Emphasis on Standardized Metrics: While standardization is a goal, there’s a potential risk that an over-emphasis on the specific metrics within EVA-MILP could inadvertently narrow the focus of instance generation research, discouraging exploration of novel instance characteristics not yet captured by the framework.
    • Mitigation: The EVA-MILP framework is designed to be extensible, and we encourage the community to propose and integrate new metrics as the field evolves.
  • Increased Computational Burden for Evaluation: thorough evaluation using EVA-MILP can be computationally intensive. This might create a barrier for researchers with limited computational resources.
    • Mitigation: Future work could explore methods for more lightweight yet informative evaluation protocols, or develop shared platforms/datasets that reduce the individual burden of running extensive evaluations.
Overall, the EVA-MILP framework is intended to be a positive contribution, aiming to improve the quality, diversity, and understanding of MILP instances used in research and development. Its primary societal impact is expected to be the indirect advancement of optimization technology, leading to better solutions for complex decision-making problems across various sectors."

### v3

@@ I.1 Methodology
K.2 Efficiency Test on Super Hard Instances
K.2.1 Experimental Objective
To validate the practical feasibility and robustness of our solver-internal feature analysis method, we designed an efficiency test. The core goal of this experiment is to demonstrate that even under strict computational constraints (i.e., tight time limits) and when facing a well-known set of hard instances, our method can still stably and efficiently extract meaningful feature distributions. This supports our key hypothesis that obtaining such deep behavioral features does not require solving instances to optimality, thereby making the evaluation both low-cost and highly efficient.
K.2.2 Experimental Methodology
Dataset Selection: We used the item_placement dataset from the ML4CO competition, which is recognized for its high solving difficulty. A total of 1000 instances were selected for testing.
Computational Budget Constraint: To simulate scenarios with limited computational resources, we imposed a strict time limit of 120 seconds for the Gurobi solver on each instance.
Validation Method: We employed a split-half validation approach to evaluate the stability of the extracted features:
  • The 1000 solver logs were randomly divided into two halves, each containing 500 instances.
  • We independently extracted solver-internal features from each half, including Root Node Gap, Heuristic Success Count, and Cut Plane Usage.
  • We then computed and compared the distributional similarity of these features across the two halves. If the random subsets originating from the same underlying distribution exhibit highly similar feature distributions (i.e., low Wasserstein distance), it indicates strong stability and reliability of the feature-extraction method.
K.2.3 Experimental Results
The split-half experiment clearly shows that even under strict time constraints, the distributions of solver-internal features extracted from the solving process remain highly stable. Key indicators and their 1-Wasserstein distances are summarized in Table 21.
Table 21: W-1 distances of key solver-internal features across two random halves
Feature Dimension W-1 Dist. Remarks
Root Node Gap 0.7613 Mean difference only 0.04%; identical standard deviation
Heuristic Success Count 0.7320 Mean and standard deviation are very close
Cut Plane Usage Compared on top 3 principal components after PCA
   – PC1 0.0917 Extremely low distance indicates highly consistent cut-plane usage pattern
   – PC2 0.0870
   – PC3 0.0904
K.2.4 Conclusion
This experiment strongly demonstrates the practical value of our framework:
  • High Efficiency and Low Cost: The results confirm that a truncated solving process (120-second timeout) is sufficient to extract stable and representative solver-internal feature distributions. This substantially reduces the time and computational cost of evaluating large-scale or extremely hard instance sets.
  • Feature Stability: The very small Wasserstein distances observed for Root Node Gap and Cut Plane Usage indicate that these early-stage solver behavior features are highly stable.
  • Practical Feasibility: The success of this test shows that the GenBench-MILP framework is not only theoretically sound but also practically applicable for evaluating MILP instances that are difficult to solve to optimality within a reasonable time frame and that reflect real-world complexity.
K.3 Tests on Other Solvers
K.3.1 Experimental Objective
To validate the generality of the GenBench-MILP framework and to investigate potential differences in solver-internal behavioral features, we applied our core solver-internal feature analysis method to two leading open-source solvers: SCIP and HiGHS. This experiment is designed to answer two key questions:
  1. 1.
    Can our framework be flexibly adapted to solvers other than Gurobi?
  2. 2.
    Do different solvers exhibit varying stability in their internal behaviors when solving the same instance set?
K.3.2 Experimental Methodology
We adopted the same split-half validation method used in the Gurobi experiments. Using the original Independent Set (IS) dataset, we randomly divided the instances into two subsets of 500 instances each. SCIP and HiGHS were then used to solve the two subsets separately, extracting three categories of solver-internal features: Root Node Gap, Heuristic Success, and Cut Plane Usage. Finally, we computed the 1-Wasserstein distance of these feature distributions between the two subsets to quantify the stability of each solver’s internal behavior. For a stable solver, the Wasserstein distance between random subsets originating from the same distribution should remain very small.
K.3.3 Experimental Results and Analysis
Our results show that, while the framework can be successfully applied to all tested solvers, their internal feature stability varies substantially.
SCIP Solver
Here are some figures 7, 8 and 9 showing the function of SCIP in spilt-half validation. For SCIP, the split-half experiment yields the following:
  • Root Node Gap: The two subsets exhibit mean gaps of 7.13% and 8.88%, respectively. The Wasserstein distance is 1.753, which is much larger than the 0.1884 observed for Gurobi, indicating lower stability for SCIP in this metric.
  • Heuristic Success: Principal Component Analysis (PCA) of six heuristic methods shows Wasserstein distances of PC1: 0.319, PC2: 0.426, and PC3: 0.778.
  • Cut Plane Usage: PCA on four cut-plane types gives Wasserstein distances of PC1: 0.700, PC2: 0.298, and PC3: 0.241.
Figure 7: Comparison of Root Node Gap distributions for SCIP in split-half validation. The two distributions represent the two random halves of the instance set.
Figure 8: Comparison of Heuristic distributions for SCIP in split-half validation. The two distributions represent the two random halves of the instance set.
Figure 9: Comparison of cut plane distributions for SCIP in split-half validation. The two distributions represent the two random halves of the instance set.
HiGHS Solver
Here are some figures 10, 11 and 12 showing the function of HiGHS in spilt-half validation. For HiGHS, the differences are even more pronounced:
  • Root Node Gap: The two subsets have mean gaps of 364.28% and 386.68%, respectively. The Wasserstein distance reaches 26.7297, indicating extremely unstable root-node relaxation behavior.
  • Heuristic Success: PCA analysis of heuristic methods gives Wasserstein distances of PC1: 0.2368 and PC2: 0.1695, with a weighted average of 0.1530, reflecting good stability.
  • Cut Plane Usage: HiGHS shows very consistent cut-plane usage. PCA reveals that the first principal component explains 100% of the variance, with a Wasserstein distance of only 0.0612, indicating highly stable cut-plane strategies.
Figure 10: Comparison of root node gap distributions for HiGHS in split-half validation. The two distributions represent the two random halves of the instance set.
Figure 11: Comparison of heuristic distributions for HiGHS in split-half validation. The two distributions represent the two random halves of the instance set.
Figure 12: Comparison of cut plane distributions for HiGHS in split-half validation. The two distributions represent the two random halves of the instance set.
K.3.4 Cross-Solver Comparison
To facilitate comparison, Table 22 summarizes key 1-Wasserstein distances from the IS split-half validation for Gurobi, SCIP, and HiGHS.
Table 22: 1-Wasserstein distances across solvers on IS split-half validation
Feature Dimension Gurobi (Baseline) SCIP HiGHS
Root Node Gap 0.1884 1.7530 26.7297
Heuristic (PC1) 0.1300 0.3190 0.2368
Cut Plane (PC1) 0.2318 0.7000 0.0612
From the table we draw three key conclusions:
  • Framework Generality: GenBench-MILP can be successfully applied to SCIP and HiGHS, confirming its potential as a general-purpose evaluation tool.
  • Stability Differences Across Solvers: The most striking observation is the difference in Root Node Gap stability. Gurobi demonstrates exceptionally high stability, SCIP shows moderate instability, and HiGHS exhibits severe instability. This implies that solver strategies (e.g., preprocessing, initial LP relaxations) may have fundamentally different sensitivities to small instance perturbations.
  • Rationale for Choosing Gurobi as Baseline: These findings strongly justify using Gurobi as the “expert evaluator” in our main study. Accurate and reliable instance-quality assessment demands a solver with highly stable internal behavior. Gurobi’s stability enables precise detection of subtle computational similarities between instances, whereas instability in other solvers can obscure these signals.
In summary, although the GenBench-MILP framework is solver-agnostic, the quality of its evaluation results is closely tied to the stability of the solver’s internal strategies. State-of-the-art commercial solvers like Gurobi provide a more precise and reliable benchmark for our evaluation framework. At the same time, these experiments reveal an additional potential application of GenBench-MILP: as a diagnostic tool to analyze and compare the internal stability and dynamic behavior of different solvers.